OpenStudy (anonymous):

A nail 3 3/4 long is driven into a board 2 1/6 inche thick. I know you need to subtract but how

OpenStudy (anonymous):

\[3\tfrac{3}{4}=\frac{15}{4}\] and \[2\tfrac{1}{6}=\frac{13}{6}\]

OpenStudy (anonymous):

use lcd of 12 to get \[\frac{15}{4}=\frac{45}{12}\] and \[\frac{13}{6}=\frac{23}{12}\] and subtract in the numeratore

OpenStudy (anonymous):

i get \[\frac{45}{12}-\frac{23}{12}=\frac{22}{12}=\frac{11}{6}\]

OpenStudy (anonymous):

except i am wrong because \[\frac{13}{6}=\frac{26}{12}\]

OpenStudy (anonymous):

so it is really \[\frac{45-26}{12}=\frac{19}{12}\]